### v2

# MaxInfo: A Training-Free Key-Frame Selection Method Using Maximum Volume for Enhanced Video Understanding

Abstract
Modern Video Large Language Models (VLLMs) often rely on uniform frame sampling for video understanding, but this approach frequently fails to capture critical information due to frame redundancy and variations in video content. We propose MaxInfo, a training-free method based on the maximum volume principle, which selects and retains the most representative frames from the input video. By maximizing the geometric volume formed by selected embeddings, MaxInfo ensures that the chosen frames cover the most informative regions of the embedding space, effectively reducing redundancy while preserving diversity. This method enhances the quality of input representations and improves long video comprehension performance across benchmarks. For instance, MaxInfo achieves a 3.28% improvement on LongVideoBench and a 6.4% improvement on EgoSchema for LLaVA-Video-7B. It also achieves a 3.47% improvement for LLaVA-Video-72B. The approach is simple to implement and works with existing VLLMs without the need for additional training, making it a practical and effective alternative to traditional uniform sampling methods.
Training-Free, Plug-and-Play, VLLMs, Video Understanding
Figure 1: Reasons for not being able to answer the correct answer at long video benchmark (Video-MME).
1 Introduction
Large language models (LLMs) such as GPT (Floridi & Chiriatti, 2020; Achiam et al., 2023), LLaMA (Touvron et al., 2023; Dubey et al., 2024), Qwen (Bai et al., 2023; Yang et al., 2024), and Mistral (Jiang et al., 2023) have revolutionized tasks like text generation, summarization, and reasoning. Recent advancements in multimodal large language models (MLLMs) (Liu et al., 2024; Goncharova et al., 2024) have extended these capabilities to include processing images, videos, and audio, enabling responses across diverse modalities. Video understanding, in particular, has garnered significant attention due to its complex, multi-dimensional nature and broad range of applications.
While models like LLaVA-Video (Zhang et al., 2024c), VideoLLaMA 2 (Cheng et al., 2024), MiniCPM-V 2.6 (Yao et al., 2024), and InternVL (Chen et al., 2024) have made strides in video understanding, they struggle with long videos due to the diversity and redundancy of video content. Uniform frame sampling, a widely used approach, often fails to capture the most informative frames, leading to missing critical details and model performance decrease. As in Figure 1 illustrate this challenge using the Video-MME (Fu et al., 2024) benchmark. As shown, uniform sampling can overlook key information essential to understanding the video, as frames critical to the answer may not be selected.
Existing approaches attempt to address these challenges by either increasing input sequence length or compressing video information. Models like LongVU (Shen et al., 2024) and MovieChat (Song et al., 2024) compress tokens per frame, while others, such as Qwen2-VL (Wang et al., 2024a) and Gemini 1.5 Pro (Team et al., 2024), process longer sequences, supporting up to 32K and 10 million tokens, respectively. However, these solutions either incur substantial computational overhead or risk losing critical temporal information. Balancing the trade-off between efficiency and accuracy remains a key challenge in long video understanding.
To address this, we propose MaxInfo – a training-free, plug-and-play method for dynamically selecting the most informative frames. Unlike uniform sampling, MaxInfo ensures that the input sequence maximizes information content and diversity. Our approach identifies and retains the most representative frames using the maximum volume principle on the matrix of frame embeddings and selects a subset of embeddings that span the most informative subspace. This ensures that redundant frames are removed while the retained frames span the most meaningful subspace of the video content.
Our contributions are as follows:
  1. 1.
    A novel framework to enhance frame diversity and informativeness. MaxInfo improves upon uniform sampling by selecting the most critical frames from a video, ensuring a more meaningful representation for VLLMs.
  2. 2.
    An advanced scene-aware extension. We extend our framework with a scene-aware algorithm that further refines frame selection by identifying key frames within individual scenes, improving performance on tasks requiring temporal coherence.
  3. 3.
    Training-free and plug-and-play integration: MaxInfo requires no retraining or fine-tuning and can be seamlessly applied to any VLLM, making it a highly practical solution for long video understanding.
Experimental results demonstrate that MaxInfo significantly improves the performance of existing models, such as LLaVA-Video (Zhang et al., 2024c), InternVL2 (Chen et al., 2024) and Qwen2-VL (Wang et al., 2024a), achieving better results on benchmarks like Video-MME (Fu et al., 2024) and EgoSchema (Mangalam et al., 2023). By overcoming the limitations of uniform sampling, MaxInfo enables more accurate and robust video understanding, ensuring that critical information from long videos is effectively captured and utilized in multimodal applications.
2 Related Works
Video Large Language Models (VLLMs). Video understanding has become a focal area of research, with numerous models excelling at video comprehension tasks. These tasks typically involve converting videos into image frames and inputting them into VLLMs. Existing approaches fall into two main categories:
Using query-based models like Q-Former (Li et al., 2023) to extract critical visual features from image frames, which are then processed by VLLMs (Li et al., 2024; Xenos et al., 2024). Encoding frame sequences with models such as CLIP (Radford et al., 2021), DINO (Caron et al., 2021), and Siglip (Zhai et al., 2023), and feeding the resulting embeddings into VLLMs (Shen et al., 2024; Jin et al., 2024a, b; Wang et al., 2024a; Chen et al., 2024; Goncharova et al., 2024). While these methods emphasize visual feature extraction and text-image semantic understanding, they often rely on uniform frame sampling or similarity-based techniques, which can overlook critical information, especially in long videos with diverse content.
Long Video Understanding. Understanding long videos poses significant challenges, primarily due to the need to balance computational efficiency with preserving critical temporal and contextual information. To address this, various strategies have been proposed:
Reducing sequence length: Methods like Video-LaVIT (Jin et al., 2024b) and LongVU (Shen et al., 2024) use cosine similarity or clustering to filter redundant frames, while MovieChat (Song et al., 2024) applies similarity thresholds for frame selection. Token compression: SlowFast-LLaVA (Xu et al., 2024) compresses visual tokens, and Chat-UniVi (Jin et al., 2024a) extracts key event tokens to reduce redundancy. Extended input lengths: Models such as Qwen2-VL (Wang et al., 2024a) and Gemini 1.5 Pro (Team et al., 2024) handle extended token lengths to process long videos, albeit with high computational costs. Despite these advancements, current methods often depend on arbitrary thresholds, fixed compression schemes, or uniform sampling, which may fail to capture the diverse and critical content of long videos effectively.
Information Maximization Techniques. Information maximization has been widely used in feature selection and dimensionality reduction across various domains. Methods like mRMR (Peng et al., 2005) and MMD (Vasconcelos, 2002) improve model performance by selecting features with high relevance and low redundancy, while MOI (Sindhwani et al., 2004) enhances classification by focusing on the most useful feature subsets. Among these, the maximum volume (MaxVol) algorithm (Goreinov et al., 2010; Mikhalev & Oseledets, 2018; Sozykin et al., 2022) stands out for its ability to maximize information by selecting linearly independent rows of a matrix, ensuring the chosen subset spans the most informative subspace. However, these methods are generally designed for static datasets and are not directly applicable to dynamic video frame selection.
In this paper, we extend the principles of MaxVol to video understanding, introducing a keyframe extraction framework tailored for VLLMs. Unlike existing approaches, our method dynamically selects diverse and representative frames. By maximizing the geometric volume of frame embeddings, MaxInfo ensures the selected frames are both diverse and highly informative. Furthermore, we advance this framework with a scene-aware algorithm, enabling finer-grained selection of critical frames for each scene. These contributions address the limitations of uniform sampling, providing a principled and effective training-free solution for long video understanding.
Figure 2: Overview of the MaxInfo Block integrated into a VLLM. We extract the most informative frames via the MaxInfo Block and then perform inference on the resulting subset of frames.
3 Method
We propose the MaxInfo Block, a plug-and-play, training-free module designed for long-video understanding tasks. It ensures both diversity of frames and comprehensive semantic coverage from a video by selecting only the most informative frames. As illustrated in Figure 2, the MaxInfo Block can be easily integrated into any VLLMs, enhancing the quality and diversity of visual information fed into the model.
3.1 Overview
Given a video, we uniformly sample $n$ frames. For example, sampling at 1 fps reduces the risk of losing important content, but still may yield a large number of frames, many of which could be redundant. This both increases computational cost and does not guarantee capturing the most informative or diverse frames. Hence, we seek a small, representative subset of frames.
Let the sampled frames be $\mathbf{I}\;=\;\{\,i_{1},\;i_{2},\;\ldots,\;i_{n}\}.$ We extract each frame’s visual representation via a CLIP-based ViT (Radford et al., 2021) and retain only the [CLS] token. Stacking these tokens yields
$\mathbf{Q}\;=\;\begin{bmatrix}q_{1}\\[4.0pt] q_{2}\\[4.0pt] \vdots\\[2.0pt] q_{n}\end{bmatrix}\;\in\;\mathbb{R}^{n\times d},$ (1)
where $q_{n}\in\mathbb{R}^{d}$ is the flattened [CLS] feature from the $n$-th frame.
3.2 Dimensionality Reduction
Handling all $n\times d$ features may still be computationally expensive when $n$ and $d$ are large. To mitigate this, we perform a truncated SVD on $\mathbf{Q}$:
$\mathbf{Q}\;=\;U\,\Sigma\,V^{T}\quad\rightarrow\quad\mathbf{Q}_{s}\;=\;U_{(:,1:s)},$ (2)
where $U\in\mathbb{R}^{n\times n}$, $\Sigma\in\mathbb{R}^{n\times d}$, and $V\in\mathbb{R}^{d\times d}$. By retaining the first $s$ singular vectors (the top $s$ columns of $U$), we obtain:
$\mathbf{Q}_{s}\;\in\;\mathbb{R}^{n\times s},$ (3)
which captures the principal visual variation among frames while drastically reducing dimensionality.
3.3 Rectangular MaxVol Frame Selection
On the next step we identify the “most informative” subset of rows of $\mathbf{Q}_{s}$, i.e., a set of frames which corresponding rows span the overall distribution of frames. To do this, we use the rectangular MaxVol algorithm (Mikhalev & Oseledets, 2018) to evaluate a submatrix of maximal volume in $\mathbf{Q}_{s}$.
For a rectangular matrix $\mathbf{A}\in\mathbb{R}^{p\times q}$, the rect-volume can be defined (up to transformations) as
[…]
$\mathbf{r}\;=\;\arg\max_{\mathbf{r}}\,\text{rect-vol}\bigl{(}\mathbf{Q}_{s}(\mathbf{r},:)\bigr{)}$ (5)
These indices $\mathbf{r}$ then specify the frames $i_{n}$ we deem most representative.
Resulting Frame Subset.
Let $r=|\mathbf{r}|$ be the number of selected frames. The final submatrix
$\mathbf{S}\;=\;\mathbf{Q}_{s}(\mathbf{r},:)$ (6)
is an $r\times s$ matrix containing the “diverse, high-information” frames. We feed only these $r$ frames into the downstream VLLM:
$\mathbf{A}_{\text{MaxInfo}}\;=\;\text{VLLM}\bigl{(}\text{Instruction},\;\mathbf{S},\;\text{Questions}\bigr{)},$ (7)
as opposed to using all $n$ frames (which might be computationally prohibitive or redundant):
$\mathbf{A}_{\text{Init}}\;=\;\text{VLLM}\bigl{(}\text{Instruction},\;\mathbf{Q},\;\text{Questions}\bigr{)}.$ (8)
Since $r\ll n$ in most cases, this yields a more efficient inference process without sacrificing essential visual context. In practice, VLLMs typically process a fixed number of frames k (e.g., every 1, 32, 64, 128 frame). Instead of directly selecting k frames or reducing the number of frames the network processes, we propose two versions of MaxInfo implementation: fast and slow. In the fast version, we apply MaxVol to the default number of frames and reduce it (e.g., reducing 256 frames to 180). In the slow version, we apply MaxVol to a larger set of frames (e.g., starting with 1024 frames and reducing it to 256), ensuring a richer pool of information for frame selection. We then apply CLIP embeddings and the MaxVol algorithm to intelligently refine this selection, maximizing the diversity and informativeness of the final k frames passed to the network. This ensures that VLLMs receive the most representative frames, improving their ability to understand long-form video content efficiently.
3.4 Fast and Slow Version
Given the differing objectives of the fast and slow versions of MaxInfo, we conducted a comparison to determine the most suitable option for each experimental setup. Additionally, we compared the two variants of MaxInfo, each designed to address different computational constraints:
  1. 1.
    Fast Version. MaxInfo is applied directly to the same number of frames as the original uniform sampling. For instance, if the base model uses $n$ frames, we retain $n$ frames and rely on MaxInfo to identify the most informative subset. This incurs minimal computational overhead and provides a quick improvement without altering the model’s default settings.
  2. 2.
    Slow Version. A larger pool of frames ($N\gg n$) is initially sampled to ensure extensive coverage. The MaxInfo Block is then applied to select the most diverse frames. If the resulting set $x$ exceeds the model’s maximum input limit $n$, we uniformly downsample $x\to n$. This approach offers potentially higher gains by starting with more frames, albeit at the cost of additional embedding computations.
3.5 Scene-Aware MaxInfo
Videos often contain multiple scenes with similar visual content per scene, making Chunk-Based MaxInfo selection suboptimal. If applied globally, MaxInfo may discard crucial frames due to artificial embedding similarities across different scenes.
To address this, we introduce Scene-Aware MaxInfo, a very simple and effective modification. Instead of performing complex scene detection, we uniformly split the video into 32 equal-sized chunks and apply MaxVol independently within each chunk. This ensures that each segment is well-represented while maintaining computational simplicity.
Chunk-Based MaxVol Selection
Given $n$ uniformly sampled frames, we split them into 32 contiguous chunks:
$\mathbf{I}=\bigcup_{i=1}^{32}\mathbf{I}^{(i)},\quad\mathbf{I}^{(i)}=\{i_{j}\mid j\in\text{chunk }i\}.$ (9)
For each chunk, we extract CLIP embeddings, apply SVD for dimension reduction, and run MaxVol to select the most representative frames.
This approach is deliberately simple but demonstrates that even Chunk-Based scene segmentation can further enhance MaxInfo’s effectiveness. It highlights the potential for more refined scene-aware selection in future work.
3.6 Summary
Our pipeline can be summarized with an Algorithm 1. Because our method is training-free and focuses purely on data selection, it can be plugged into diverse video-focused VLLMs to reduce cost and improve performance on long-video tasks.
Algorithm 1 MaxInfo Block: SVD + MaxVol for Keyframe Selection
1:  Input: A set of $n$ frames $\mathbf{I}=\{i_{1},i_{2},\dots,i_{n}\}$
2:  Embedding: Convert each frame $i_{j}$ into a [CLS] embedding:
$q_{n}=\mathrm{flatten}\bigl{(}\mathrm{clip\_model}(i_{n})\bigr{)},\mathbf{Q}=\begin{bmatrix}q_{1}\\ q_{2}\\ \vdots\\ q_{n}\end{bmatrix}\in\mathbb{R}^{n\times d}.$
3:  SVD Reduction: Perform truncated SVD on $\mathbf{Q}$:
$\mathbf{Q}\approx\mathbf{U}_{r}\,\mathbf{\Sigma}_{r}\,\mathbf{V}_{r}^{\top}\quad\rightarrow\quad\mathbf{Q_{s}}=\mathbf{U}_{r}\in\mathbb{R}^{n\times r}.$
[…]
$\mathrm{piv}=\text{rect\_maxvol}(\mathbf{Q_{s}},\mathrm{Tol}),$
identifying rows (frames) that span the reduced embedding space.
5:  Output: Indices $\mathrm{piv}$ of the most informative keyframes.
4 Experiments
4.1 Implementation Details
In order to evaluate the contribution of MaxInfo to video understanding, we employed widely-used video understanding benchmarks, covering short-video tasks (such as MVBench (Li et al., 2024)) and medium-to-long video tasks (such as EgoSchema (Mangalam et al., 2023), Video-MME (Fu et al., 2024), and LongVideoBench (Shen et al., 2024)). For complete fairness, we only compared improved versions of the model against itself without MaxVol with freezing generation parameters, seed and prompt. We mostly focus on longer videos because better frame selection plays a bigger role in longer, more complex videos, whereas shorter ones intuitively work well with uniform sampling due to their lower information content and complexity.
We ensured that the resulting sequence length of a set of visual and textual tokens did not exceed the maximum sequence length for this LLM. When evaluating models using MaxInfo, we limited the number of selected frames so that they did not exceed the maximum allowed for the context of the estimated VLLM. For the evaluation on all benchmarks, we have set the generation temperature to 0.
For the general multiple-choice question-answering evaluation, we follow the official guidelines to construct the instructions using the provided questions and options. We added a prompt to the question and options like ”Respond with only the letter (A, B, C, or D) of the correct option.” for LongVideoBench (Wu et al., 2024), Video-MME (Fu et al., 2024), MLVU (Zhou et al., 2024) and MVBench (Li et al., 2024) or ”Answer with the option’s letter from the given choices directly and only give the best option.” for EgoSchema (Mangalam et al., 2023). We follow the original benchmarks setup to calculate the final scores, and we also align our evaluation protocols with other evaluation toolkits, such as lmms-eval (Zhang et al., 2024a).
To ensure the reproducibility of our results, we have included the main hyperparameters used for all benchmarks and estimated models in the results tables, such as tolerance and rank for the MaxInfo algorithm, the number of sampled frames, and the number of initial frames (before MaxInfo).
4.2 Main Results
Overall Performance. Table 1, Table 2 and Table 3 present the performance gains achieved by integrating MaxInfo into existing InternVL (Chen et al., 2024), Qwen2-VL (Wang et al., 2024a) and LlaVA-Video (Zhang et al., 2024c) models. The experimental results show that MaxInfo Block exhibits significant performance improvement on a wide range of models. In particular, in the LLaVA-Video-7B and Qwen-VL-2B models, the introduction of MaxInfo Block improves the accuracy by 0.9%/1.7%, 6.4%, 3.3% and 1.4%/1.2%, 2.3%, 1.5% in VideoMME, EgoSchema, and LongVideoBench, respectively, which is significantly better than the versions without the block. This improvement not only validates the effectiveness of MaxInfo Block, but also provides new ideas for future research on video understanding tasks.
Table 1: Comparing Video-LLMs with and without MaxInfo as an additional module. Except for ours (-MaxInfo) and the models with *, all results are copied from the related papers of benchmarks or models. Base models uses uniform frames sampling. For the Video-MME benchmark, we report results under two standard settings: without subtitles (wo. sub.) and with subtitles (w. sub.). The % accuracy sign is omitted for clarity.
Model Size VideoMME (wo/w-subs) Egoshcema LongVideoBench
LLaVA-Video (Zhang et al., 2024c) 7B 63.3/69.7(64) 57.3(64) 58.2(64)
+ MaxInfo 7B 64.2/71.464→(6,64) 63.7128→(64,64) 61.5128→(1,64)
$\triangle$ +0.9%/+1.7% +6.4% +3.3%
LLaVA- (Zhang et al., 2024c) 72B 70.5/76.9(64) 65.6(64) 61.9(64)
+ MaxInfo 72B 70.2/77.664→(6,64) 69.4128→(64,64) 64.9128→(64,1)
$\triangle$ -0.3%/+0.7% +3.8% +3%
Qwen2-VL (Wang et al., 2024a) 2B 55.6/60.4(786) 54.9(180) 47.3(256)
+ MaxInfo 2B 57.0/61.6256→(254,4) 57.2180→(180,12) 48.8256→(224,1)
$\triangle$ +1.4%/+1.2% +2.3% +1.5%
Qwen2-VL (Wang et al., 2024a) 7B 63.3/69.0(768) 66.7(180) 53.7(256)
+ MaxInfo 7B 62.1/70.0256→(254,4) 64.3180→(180,12) 55.7256→(224,1)
$\triangle$ -1.2/+1.0% +2.4% +2.0%
Table 2: Comparison of the InternVL2 model series with and without MaxInfo as an additional module. InternVL2 uses uniform frame sampling and has a limited context size. Therefore, the maximum number of frames for evaluation across all benchmarks was limited to 32. The maximum number of input and output frames for the MaxInfo module is set to 128 and 32 respectively. For the LongVideoBench benchmark, we report results under a default setting for video LLMs, i.e. without subtitles. The % accuracy sign is omitted for clarity.
Model Size MVBench Egoschema LongVideoBench
InternVL2 (Chen et al., 2024) 1B 55.42 34.02 43.38
+ MaxInfo(128,32) 1B 56.07 34.11 42.78
$\triangle$ +0.65% +0.09% -0.6%
InternVL2 (Chen et al., 2024) 2B 59.72 46.10 43.97
+ MaxInfo(128,32) 2B 59.02 46.33 47.27
$\triangle$ -0.7% +0.23% +3.3%
Table 3: Performance comparison between MaxInfo Block and other key frame selection methods on the Video-MME-Long benchmark. Results are measured without adding subtitles. $Q$ refers to the Qwen2-VL-7B model and $L$ refers to the LLaVA-Video-7B model. The % accuracy sign is omitted for clarity.
Model Size frames Acc.
Video-Tree (Wang et al., 2024c) - - 54.2/-
Long-VU (Shen et al., 2024) 7B 1fps -/60.6
Chat-Univi (Jin et al., 2024a) 7B 64 35.8/-
Q+MaxInfo 7B 256 $\rightarrow$ (254, 4) 51.8/63.9
L+MaxInfo 7B 64 $\rightarrow$ (64, 6) 54.6/65.9
MLVU Benchmark. Table 4 evaluates both MaxInfo and the benefits of Chunk-Based Scene-Awareness on the MLVU benchmark, demonstrating consistent improvements over uniform sampling. While these improvements are modest, it is important to highlight that the method is exceptionally simple and entirely training-free. This simplicity underscores the potential of even basic scene-awareness to enhance performance, suggesting that more sophisticated scene detection techniques could lead to further gains. Such results indicate that scene-awareness is a promising direction for improving video understanding tasks.
Table 4: Comparison of performance on the MLVU benchmark between the original versions of the models in the Qwen-VL family and their variants with the MaxInfo and Naive Scene-Aware (NSA) additional modules. The original models use uniform sampling with a maximum of 128 frames. The additional modules use from 128 to 1024 frames per input. The best results are highlighed in bold, second best are underlined.
Model Size Acc.
Qwen2-VL (Wang et al., 2024a) 2B 52.18
+ MaxVol 2B 52.37
+ Chunk-Based 2B 52.69
Qwen2-VL (Wang et al., 2024a) 7B 64.32
+ MaxVol 7B 64.59
+ Chunk-Based 7B 64.82
4.3 Performance Analysis: MaxInfo vs. Uniform Sampling
To better understand the strengths and trade-offs of MaxInfo, we analyzed per-task accuracy across multiple benchmarks.
Figure 3: Accuracy comparison between Uniform Sampling and MaxInfo on Video-MME (Qwen2-VL-2B).
Figure 4: Accuracy comparison between Uniform Sampling and MaxInfo on LongVideoBench (Qwen2-VL-2B).
Figure 5: Accuracy comparison between Uniform Sampling and MaxInfo on MLVU (Qwen2-VL-7B) with 64 final and 1024 initial frames.
Our results, as shown in Figure 3, Figure 4, and Figure 5, indicate that MaxInfo outperforms uniform sampling on tasks requiring high information density (e.g., counting, summarization, and spatial reasoning). However, for tasks requiring fine-grained temporal reasoning (e.g., object recognition and action perception), uniform sampling retains a slight advantage. These findings highlight a fundamental trade-off between information maximization and temporal continuity, suggesting potential directions for hybrid selection methods.
4.4 Fast vs. Slow Version Comparisons
As Table 5 shows, the slow version outperforms the fast version in most cases, particularly when processing longer videos. The initial oversampling mechanism in the slow version provides MaxInfo with a richer selection space, which significantly improves performance. However, experimental results indicate that the fast version may outperform the slow version in specific cases, depending on the benchmark type and the number of initial samples. Nevertheless, in scenarios with limited computational resources or strict latency requirements, the fast version remains a practical and reliable choice due to its efficiency. This flexibility allows the two versions to adapt to different task requirements, offering users more options.
Table 5: Comparison of the impact of the slow and fast versions of the MaxInfo module on the performance of the LLaVA-Video-7B model on the LongVideoBench. All results are reproduced, except those marked with *, which are copied from the corresponding papers (Zhang et al., 2024c).
Model Input max. frames Output frames range Avg. frames Acc.
InternVL2 1B (Chen et al., 2024) 32 - 32 43.38
+ MaxInfofast 32 (1, 32) 30 43.6 +0.22%
+ MaxInfoslow 64 (1, 16) 16 44.65 +1.05%
InternVL2 2B (Chen et al., 2024) 32 - 32 46.97
+ MaxInfofast 32 (1, 32) 30 47.27 +0.73%
+ MaxInfoslow 64 (1, 16) 16 46.82 -0.15%
LLaVA-Video 7B (Zhang et al., 2024c) 64 - 64 58.2
+ MaxInfofast 64 (1, 64) 52 60.21 +2.01%
+ MaxInfoslow 128 (1, 64) 58 61.48 +3.28%
LLaVA-Video 72B (Zhang et al., 2024c) 64 - 64 61.9
+ MaxInfofast 64 (1, 64) 52 65.37 +3.47%
+ MaxInfoslow 128 (1, 64) 58 64.92 +3.02%
4.5 Ablation Study
To further evaluate the effectiveness of MaxInfo, we conducted a series of ablation experiments focusing on two key aspects: the impact of the choice of visual encoder and the influence of key hyperparameters in the MaxInfo module, particularly tolerance and rank.
4.5.1 Vision Encoder Impact
Table 6 presents an ablation study on different visual encoders used for frame representation in MaxInfo.
Table 6: An ablation study on the choice of a visual encoder to obtain the visual representation of the frames in the MaxInfo approach. The best results are highlighed in bold, second best are underlined. Based on LLaVA-Video-7B (Zhang et al., 2024c)
Model Size Visual encoder Param. Acc.
LLava-Video 7B CLIP-ViT-Large 427.9 M. 58.94
LLava-Video 7B CLIP-ViT-Base 149.6 M. 58.79
LLava-Video 7B DINO2-base 86.6 M. 58.94
LLava-Video 7B DINO2-large 304.4 M. 58.86
LLava-Video 7B SigLIP-base-224 203.2 M. 59.76
LLava-Video 7B SigLIP-base-384 878 M. 59.24
We evaluated CLIP, DINOv2, and SigLIP as visual encoders. DINOv2 performed comparably to CLIP, despite lacking vision-language alignment. However, SigLIP outperformed both CLIP and DINOv2, likely due to its stronger language-vision connectivity. Interestingly, the larger SigLIP model underperformed, suggesting that more complex encoders require careful hyperparameter tuning (e.g., tolerance and rank) for optimal frame selection.
4.5.2 Impact of MaxInfo Block Hyperparameters
This section examines the effect of MaxInfo parameters (rank R and tolerance Tol) as well as the initial number of samples on the final model accuracy.
Effect of Different Tolerances and Ranks on the Model with Fixed Sampling. We conducted a series of tests with fixed benchmarks, model settings, and an initial pool of frames (e.g., $n^{*}=96$). As shown in Figure 6, the best observed result achieved a 3.3% improvement over the base LLaVA-Video-7B model with $\text{Tol}=0.3$ and $\text{R}=8$.
Figure 6: Effect of initial sampling on MaxInfo performance for LlaVa-Video 7B model (fixed R=8, Tol=0.15).
From these experiments, we derived the following guidelines:
  1. 1.
    Large Rank, Moderate Tol: When the Rank $R$ is large, a moderate Tolerance (Tol) helps avoid selecting overly similar frames, optimizing the trade-off between diversity and redundancy.
  2. 2.
    Balanced Selection: An excessively large $R$ combined with a very small Tol may cause the frame set to grow excessively, potentially exceeding the model’s capacity. Thus, $R$ and Tol should be selected based on the maximum feasible input frame count for a given LLM.
  3. 3.
    Performance Sensitivity: Our experiments show that performance is most sensitive to Tol values in the range $\text{Tol}\in[0.1,0.6]$. Beyond this range, improvements plateau or regress due to over-pruning or under-pruning.
  4. 4.
    Convergence: When setting $R=12$ and varying Tol values to 0.30, 0.45, and 0.60, the best result converged at 61.1%.
Effect of Sampling on Accuracy (Fixed Rank and Tolerance). We analyzed how varying the initial number of sampled frames impacts accuracy, while keeping the hyperparameters fixed (R=8, Tol=0.15). As shown in Figure 7, accuracy initially improves with the addition of more frames, but beyond a certain threshold, it begins to plateau or slightly decline. Our key observations are as follows:
Figure 7: Comparing the effect of initial sampling on the MaxInfo Block, n represents the initial number of frames sampled, and after passing through the MaxInfo Block, the number of frames is increased to a maximum of 64.
  1. 1.
    Increasing the initial number of frames provides more diverse information, enabling MaxVol to select better keyframes.
  2. 2.
    Beyond a certain number of frames, accuracy slightly declines due to excessive redundancy, which hampers MaxVol’s efficiency in keyframe selection.
  3. 3.
    There exists an optimal trade-off between the initial frame count and computational cost. In our tests, 128 as initial frames yielded the best accuracy for the LLaVA-Video-7B model.
5 Conclusion
In this work, we introduced MaxInfo – a training-free method for selecting the most informative frames from videos, improving VLLMs inference. Our results consistently demonstrate that informative frame selection outperforms uniform sampling, leading to improved performance of state of the art VLLMs (LLaVA-Video, InternVL and Qwen2-VL with different sizes) across multiple benchmarks. However, the improvements are relatively small, which is expected given the training-free nature and the simplicity of our approach.
Beyond demonstrating empirical gains, we believe our work will encourage the community to focus more on frame selection strategies, an often-overlooked aspect of video understanding. Additionally, we have shown that even minimal refinements, such as chunk-wise MaxVol in our Scene-Aware MaxInfo, can further enhance results, demonstrating that simple adjustments can lead to meaningful improvements.
Finally, we hypothesize that training VLLMs with informative frame sampling, rather than simple uniform frame selection, could further enhance their capabilities when later used with inference-time MaxInfo techniques. We hope this work serves as a foundation for future research into more efficient, information-aware video sampling strategies for large-scale multimodal learning.
Impact Statement
This work introduces a training-free framework for improving frame sampling in Vision-Language Large Models (VLLMs), enhancing video understanding tasks. Such advancements are crucial for applications in education, accessibility, and public safety, where accurate and efficient video analysis can have a meaningful impact.
However, improved video analysis capabilities may also raise ethical concerns, including potential misuse in surveillance, privacy violations, or biases in automated systems. Ensuring responsible deployment with a focus on fairness and transparency is essential to mitigate these risks.
In summary, while our approach offers significant benefits, its adoption must adhere to ethical principles to promote equitable and responsible use.
Appendix A Appendix
A.1 Visualisation of MaxInfo Block results
Figure 8: MaxInfo Block sampled frames.
As shown Figure 8, MaxInfo Block can help us to reduce the redundancy of video information, better understand the video content and improve the model performance by selecting the most informative and representative key frames among multiple frames.
A.2 Additional comparisons with other models
We compared the performance of VLLMs of the same 7B size, as shown in Table 7, such as Open Source Model, Proprietary Models and Train-Free Model or Key Frames selection, on Video-MME Benchmark.
Table 7: Comparison of model results on Video-MME Benchmark.
Model LLM Size Frames Overall (WO/W-Subs) Short (WO/W-Subs) Medium (WO/W-Subs) Long (WO/W-Subs)
Proprietary Models
GPT-4o (OpenAI, 2024) - $384^{2*}$ 71.2/77.8 80.1/82.2 71.3/76.8 62.2/74.3
Gemini 1.5 Pro (Team et al., 2024) - 1/0.5 fps 75.0/81.3 81.7/84.5 74.3/81.0 67.4/77.4
Open Source Model
Video-LLaVA (Lin et al., 2023) 7B 8 39.9/41.6 45.3/46.1 38.0/40.7 36.2/38.1
Long-LLaVA (Wang et al., 2024b) 7B 64 52.9/57.1 61.9/66.2 51.4/54.7 45.4/50.3
mPLUG-Owl3 (Ye et al., 2024) 7B 128 59.3/68.1 70.0/72.8 57.7/66.9 50.1/64.5
LongVA (Zhang et al., 2024b) 7B 128 52.6/54.3 61.1/61.1 50.4/53.6 46.2/47.6
LongVILA (Xue et al., 2024) 7B 128 57.5/60.6 69.3/70.4 56.1/59.2 47.0/52.1
256 57.4/61.8 69.0/70.8 56.7/60.6 46.4/54.0
Train-Free Model or Key Frames selection
Video-Tree* (Wang et al., 2024c) - - -/- -/- -/- 54.2/-
LongVU* (Shen et al., 2024) 7B 1fps -/- -/- -/- -/60.6
Chat-UniVi (Jin et al., 2024a) 7B 64 40.6/- 45.7/- 40.3/- 35.8/-
MaxInfo Method
Qwen2-VL-MaxInfo 7B Avg.180 62.1/70.0 72.5/76.4 62.0/69.8 51.8/63.9
LLaVA-Video-MaxInfo 7B Avg.58 64.2/71.4 74.6/78.7 63.3/69.8 54.6/65.9
A.3 Impact of MaxInfo Block’s Tol and Rank on Models.
Overview of the details of how Tol and R affect the accuracy of the model, as Table 8, I.Frame is the number of frames input to the MaxInfo module, and F.Frame is the maximum number of frames after going through the MaxInfo Block. The effect of the initial number of samples on the model as shown in Table 9.
Table 8: Comparison of experimental results on the effect of different tol and rank on the LLava-Video (Zhang et al., 2024c).
Model Size R Tol (I.Frames, F.Frames) Acc.
LLava-Video 7B 12 0.15 (96, 64) 60.35
LLava-Video 7B 12 0.30 (96, 64) 61.25
LLava-Video 7B 12 0.45 (96, 64) 61.10
LLava-Video 7B 12 0.60 (96, 64) 61.14
LLava-Video 7B 8 0.15 (96, 64) 60.06
LLava-Video 7B 8 0.30 (96, 64) 61.48
LLava-Video 7B 8 0.45 (96, 64) 59.68
LLava-Video 7B 8 0.60 (96, 64) 57.80
LLava-Video 7B 6 0.10 (96, 64) 61.10
LLava-Video 7B 6 0.15 (96, 64) 60.96
LLava-Video 7B 6 0.30 (96, 64) 60.58
LLava-Video 7B 6 0.45 (96, 64) 59.01
LLava-Video 7B 6 0.60 (96, 64) 57.89
Table 9: Comparison of the effect of different initial sampling values on the model.
Model Size (I.Frames, F.Frames) Acc.
LLava-Video 7B (256, 64) 60.58
LLava-Video 7B (192, 64) 60.73
LLava-Video 7B (128, 64) 60.88
LLava-Video 7B (96, 64) 60.06

### v3

@@
 Modern Video Large Language Models (VLLMs) often rely on uniform frame sampling for video understanding, but this approach frequently fails to capture critical information due to frame redundancy and variations in video content.
-We propose MaxInfo, a training-free method based on the maximum volume principle, which selects and retains the most representative frames from the input video.
+We propose MaxInfo, the first training-free method based on the maximum volume principle, which is available in Fast and Slow versions and a Chunk-based version that selects and retains the most representative frames from a video.
 By maximizing the geometric volume formed by selected embeddings, MaxInfo ensures that the chosen frames cover the most informative regions of the embedding space, effectively reducing redundancy while preserving diversity.
@@
 For instance, MaxInfo achieves a 3.28% improvement on LongVideoBench and a 6.4% improvement on EgoSchema for LLaVA-Video-7B.
-It also achieves a 3.47% improvement for LLaVA-Video-72B.
-The approach is simple to implement and works with existing VLLMs without the need for additional training, making it a practical and effective alternative to traditional uniform sampling methods.
-Training-Free, Plug-and-Play, VLLMs, Video Understanding
-Figure 1: Reasons for not being able to answer the correct answer at long video benchmark (Video-MME).
+Moreover, MaxInfo boosts LongVideoBench performance by 3.47% on LLaVA-Video-72B and 3.44% on MiniCPM4.5.
+The approach is simple to implement and works with existing VLLMs without the need for additional training and very lower latency, making it a practical and effective alternative to traditional uniform sampling methods.
+Our code are available at https://github.com/FusionBrainLab/MaxInfo.git
+Figure 1: Reasons why the Uniform Sampling approach cannot answer the correct answer in long videos. An example of MaxInfo’s sampling approach.
 1 Introduction
-Large language models (LLMs) such as GPT (Floridi & Chiriatti, 2020; Achiam et al., 2023), LLaMA (Touvron et al., 2023; Dubey et al., 2024), Qwen (Bai et al., 2023; Yang et al., 2024), and Mistral (Jiang et al., 2023) have revolutionized tasks like text generation, summarization, and reasoning.
-Recent advancements in multimodal large language models (MLLMs) (Liu et al., 2024; Goncharova et al., 2024) have extended these capabilities to include processing images, videos, and audio, enabling responses across diverse modalities.
+Large language models (LLMs) such as GPT [gpt-3, gpt4], LLaMA [llama2, llama3], Qwen [qwen, qwen2], and Mistral [jiang2023mistral] have revolutionized tasks like text generation, summarization, and reasoning.
+Recent advancements in multimodal large language models (MLLMs) [llava, omnifusion] have extended these capabilities to include processing images, videos, and audio, enabling responses across diverse modalities.
 Video understanding, in particular, has garnered significant attention due to its complex, multi-dimensional nature and broad range of applications.
-While models like LLaVA-Video (Zhang et al., 2024c), VideoLLaMA 2 (Cheng et al., 2024), MiniCPM-V 2.6 (Yao et al., 2024), and InternVL (Chen et al., 2024) have made strides in video understanding, they struggle with long videos due to the diversity and redundancy of video content.
+While models like LLaVA-Video [llava-video], VideoLLaMA 2 [cheng2024videollama2], MiniCPM-V 2.6 [yao2024minicpmv], and InternVL [internvl] have made strides in video understanding, they struggle with long videos due to the diversity and redundancy of video content.
 Uniform frame sampling, a widely used approach, often fails to capture the most informative frames, leading to missing critical details and model performance decrease.
-As in Figure 1 illustrate this challenge using the Video-MME (Fu et al., 2024) benchmark.
+As shown in Figure 1 illustrate this challenge using the Video-MME [video-mme] benchmark.
 As shown, uniform sampling can overlook key information essential to understanding the video, as frames critical to the answer may not be selected.
 Existing approaches attempt to address these challenges by either increasing input sequence length or compressing video information.
-Models like LongVU (Shen et al., 2024) and MovieChat (Song et al., 2024) compress tokens per frame, while others, such as Qwen2-VL (Wang et al., 2024a) and Gemini 1.5 Pro (Team et al., 2024), process longer sequences, supporting up to 32K and 10 million tokens, respectively.
+Models like LongVU [longvu] and MovieChat [song2024moviechat] compress tokens per frame, while others, such as Qwen2-VL [qwen2-vl] and Gemini 1.5 Pro [team2024gemini], process longer sequences, supporting up to 32K and 10 million tokens, respectively.
 However, these solutions either incur substantial computational overhead or risk losing critical temporal information.
@@
     Training-free and plug-and-play integration: MaxInfo requires no retraining or fine-tuning and can be seamlessly applied to any VLLM, making it a highly practical solution for long video understanding.
-Experimental results demonstrate that MaxInfo significantly improves the performance of existing models, such as LLaVA-Video (Zhang et al., 2024c), InternVL2 (Chen et al., 2024) and Qwen2-VL (Wang et al., 2024a), achieving better results on benchmarks like Video-MME (Fu et al., 2024) and EgoSchema (Mangalam et al., 2023).
-By overcoming the limitations of uniform sampling, MaxInfo enables more accurate and robust video understanding, ensuring that critical information from long videos is effectively captured and utilized in multimodal applications.
 2 Related Works
@@
 These tasks typically involve converting videos into image frames and inputting them into VLLMs.
-Existing approaches fall into two main categories:
-Using query-based models like Q-Former (Li et al., 2023) to extract critical visual features from image frames, which are then processed by VLLMs (Li et al., 2024; Xenos et al., 2024).
-Encoding frame sequences with models such as CLIP (Radford et al., 2021), DINO (Caron et al., 2021), and Siglip (Zhai et al., 2023), and feeding the resulting embeddings into VLLMs (Shen et al., 2024; Jin et al., 2024a, b; Wang et al., 2024a; Chen et al., 2024; Goncharova et al., 2024).
+Existing approaches fall into two main categories: Using query-based models like Q-Former [blip2] to extract critical visual features from image frames, which are then processed by VLLMs [mvbench, xenos2024vllms].
+Encoding frame sequences with models such as CLIP [clip], DINO [dino], and Siglip [Siglip], and feeding the resulting embeddings into VLLMs [longvu, Chat-univi, video-lavit, qwen2-vl, internvl, omnifusion, mvbench, chen2024sharegpt4video].
 While these methods emphasize visual feature extraction and text-image semantic understanding, they often rely on uniform frame sampling or similarity-based techniques, which can overlook critical information, especially in long videos with diverse content.
@@
 Understanding long videos poses significant challenges, primarily due to the need to balance computational efficiency with preserving critical temporal and contextual information.
-To address this, various strategies have been proposed:
-Reducing sequence length: Methods like Video-LaVIT (Jin et al., 2024b) and LongVU (Shen et al., 2024) use cosine similarity or clustering to filter redundant frames, while MovieChat (Song et al., 2024) applies similarity thresholds for frame selection.
-Token compression: SlowFast-LLaVA (Xu et al., 2024) compresses visual tokens, and Chat-UniVi (Jin et al., 2024a) extracts key event tokens to reduce redundancy.
-Extended input lengths: Models such as Qwen2-VL (Wang et al., 2024a) and Gemini 1.5 Pro (Team et al., 2024) handle extended token lengths to process long videos, albeit with high computational costs.
+To address this, various strategies have been proposed, Reducing sequence length: Methods like Video-LaVIT [video-lavit] and LongVU [longvu] use cosine similarity or clustering to filter redundant frames, while MovieChat [song2024moviechat] applies similarity thresholds for frame selection.
+Token compression: SlowFast-LLaVA [slowfast] compresses visual tokens, and Chat-UniVi [Chat-univi] extracts key event tokens to reduce redundancy.
+Extended input lengths: Models such as Qwen2-VL [qwen2-vl] and Gemini 1.5 Pro [team2024gemini] handle extended token lengths to process long videos, albeit with high computational costs.
+In addition, numerous keyframe extraction algorithms have been proposed, such as LongVA [zhang2024long], Frame-Voyager [yu2024frame], AKS [tang2025adaptive], VideoTree [videotree], M-LLM [hu2025m], BOLT [liu2025bolt], VSLS [guo2025logic], AdaReTaKe [wang2025adaretake], Q-Frame [zhang2025q], GenS [yao2025generative], and ViLaMP [cheng2025scaling], all of which have demonstrated remarkable performance across various long video understanding tasks.
 Despite these advancements, current methods often depend on arbitrary thresholds, fixed compression schemes, or uniform sampling, which may fail to capture the diverse and critical content of long videos effectively.
 Information Maximization Techniques.
-Information maximization has been widely used in feature selection and dimensionality reduction across various domains.
-Methods like mRMR (Peng et al., 2005) and MMD (Vasconcelos, 2002) improve model performance by selecting features with high relevance and low redundancy, while MOI (Sindhwani et al., 2004) enhances classification by focusing on the most useful feature subsets.
-Among these, the maximum volume (MaxVol) algorithm (Goreinov et al., 2010; Mikhalev & Oseledets, 2018; Sozykin et al., 2022) stands out for its ability to maximize information by selecting linearly independent rows of a matrix, ensuring the chosen subset spans the most informative subspace.
-However, these methods are generally designed for static datasets and are not directly applicable to dynamic video frame selection.
-In this paper, we extend the principles of MaxVol to video understanding, introducing a keyframe extraction framework tailored for VLLMs.
-Unlike existing approaches, our method dynamically selects diverse and representative frames.
-By maximizing the geometric volume of frame embeddings, MaxInfo ensures the selected frames are both diverse and highly informative.
-Furthermore, we advance this framework with a scene-aware algorithm, enabling finer-grained selection of critical frames for each scene.
-These contributions address the limitations of uniform sampling, providing a principled and effective training-free solution for long video understanding.
-Figure 2: Overview of the MaxInfo Block integrated into a VLLM. We extract the most informative frames via the MaxInfo Block and then perform inference on the resulting subset of frames.
+Information maximization is widely used for feature selection and dimensionality reduction.
+Methods such as mRMR [peng2005feature] and MMD [MMD] improve model performance by selecting features with high relevance and low redundancy, while MOI [MOI] focuses on the most informative feature subsets to enhance classification.
+The maximum volume (MaxVol) algorithm [goreinov2010find, Rectmax, sozykin2022ttopt] selects linearly independent rows of a matrix to cover the most informative subspace.
+In this paper, we extends the MaxVol principle to video understanding and proposes a keyframe extraction framework tailored for VLLMs.
+Unlike existing methods, our approach dynamically selects diverse and representative frames.
+By maximizing the geometric volume of frame embeddings, MaxInfo ensures that the selected frames are both informative and diverse, and, combined with a scene-aware algorithm, enables fine-grained selection of keyframes for each video segment, providing an efficient, training-free solution for long video understanding.
 3 Method
@@
 As illustrated in Figure 2, the MaxInfo Block can be easily integrated into any VLLMs, enhancing the quality and diversity of visual information fed into the model.
+Figure 2: Overview of the MaxInfo Block integrated into a VLLM. We extract the most informative frames via the MaxInfo Block and then perform inference on the resulting subset of frames.
 3.1 Overview
@@
 Hence, we seek a small, representative subset of frames.
-Let the sampled frames be $\mathbf{I}\;=\;\{\,i_{1},\;i_{2},\;\ldots,\;i_{n}\}.$ We extract each frame’s visual representation via a CLIP-based ViT (Radford et al., 2021) and retain only the [CLS] token.
+Let the sampled frames be $\mathbf{I}\;=\;\{\,i_{1},\;i_{2},\;\ldots,\;i_{n}\}.$ We extract each frame’s visual representation via a CLIP-based ViT [clip] and retain only the [CLS] token.
 Stacking these tokens yields
-$\mathbf{Q}\;=\;\begin{bmatrix}q_{1}\\[4.0pt] q_{2}\\[4.0pt] \vdots\\[2.0pt] q_{n}\end{bmatrix}\;\in\;\mathbb{R}^{n\times d},$ (1)
+$\mathbf{Q}=\begin{bmatrix}q_{1}&q_{2}&\cdots&q_{n}\end{bmatrix}^{\top}$ (1)
 where $q_{n}\in\mathbb{R}^{d}$ is the flattened [CLS] feature from the $n$-th frame.
@@
 On the next step we identify the “most informative” subset of rows of $\mathbf{Q}_{s}$, i.e., a set of frames which corresponding rows span the overall distribution of frames.
-To do this, we use the rectangular MaxVol algorithm (Mikhalev & Oseledets, 2018) to evaluate a submatrix of maximal volume in $\mathbf{Q}_{s}$.
+To do this, we use the rectangular MaxVol algorithm [Rectmax] to evaluate a submatrix of maximal volume in $\mathbf{Q}_{s}$.
 For a rectangular matrix $\mathbf{A}\in\mathbb{R}^{p\times q}$, the rect-volume can be defined (up to transformations) as
-[…]
-[…]
+$\text{rect-vol}(\mathbf{A})\;=\;\sqrt{\det(\mathbf{A}\,\mathbf{A}^{T})}$ (4)
+Maximizing $\text{rect-vol}(\mathbf{A})$ with respect to the selection of rows corresponds to identifying the subset of frames that best preserves the variation in the data.
+We denote the selected row indices as
 $\mathbf{r}\;=\;\arg\max_{\mathbf{r}}\,\text{rect-vol}\bigl(\mathbf{Q}_{s}(\mathbf{r},:)\bigr)$ (5)
-These indices $\mathbf{r}$ then specify the frames $i_{n}$ we deem most representative.
+where indices $\mathbf{r}$ then specify the frames $i_{n}$ we deem most representative.
 Resulting Frame Subset.
 Let $r=|\mathbf{r}|$ be the number of selected frames. The final submatrix
-$\mathbf{S}\;=\;\mathbf{Q}_{s}(\mathbf{r},:)$ (6)
-is an $r\times s$ matrix containing the “diverse, high-information” frames. We feed only these $r$ frames into the downstream VLLM:
+$\mathbf{S}\;=\;\mathbf{Q}(\mathbf{r},:)$ (6)
+is an $r\times n$ matrix containing the diverse, high-information frames. We feed only these $r$ frames into the downstream VLLM:
 $\mathbf{A}_{\text{MaxInfo}}\;=\;\text{VLLM}\bigl(\text{Instruction},\;\mathbf{S},\;\text{Questions}\bigr),$ (7)
@@
 $\mathbf{A}_{\text{Init}}\;=\;\text{VLLM}\bigl(\text{Instruction},\;\mathbf{Q},\;\text{Questions}\bigr).$ (8)
-Since $r\ll n$ in most cases, this yields a more efficient inference process without sacrificing essential visual context.
-In practice, VLLMs typically process a fixed number of frames k (e.g., every 1, 32, 64, 128 frame).
-Instead of directly selecting k frames or reducing the number of frames the network processes, we propose two versions of MaxInfo implementation: fast and slow.
-In the fast version, we apply MaxVol to the default number of frames and reduce it (e.g., reducing 256 frames to 180).
-In the slow version, we apply MaxVol to a larger set of frames (e.g., starting with 1024 frames and reducing it to 256), ensuring a richer pool of information for frame selection.
-We then apply CLIP embeddings and the MaxVol algorithm to intelligently refine this selection, maximizing the diversity and informativeness of the final k frames passed to the network.
-This ensures that VLLMs receive the most representative frames, improving their ability to understand long-form video content efficiently.
+Given that the value of r is much smaller than n (i.e., $r\ll n$) in most application scenarios, the design significantly improves the inference efficiency while effectively avoiding the loss of critical visual context information.
+We innovatively propose the MaxInfo algorithm framework scene-aware MaxInfo and fast and slow versions of the implementation scheme.
+Experimental results show that both exhibit significant enhancements.
 3.4 Fast and Slow Version
@@
 Additionally, we compared the two variants of MaxInfo, each designed to address different computational constraints:
-  1. 1.
 Fast Version.
@@
 This incurs minimal computational overhead and provides a quick improvement without altering the model’s default settings.
-  2. 2.
 Slow Version.
@@
 This approach offers potentially higher gains by starting with more frames, albeit at the cost of additional embedding computations.
-3.5 Scene-Aware MaxInfo
-Videos often contain multiple scenes with similar visual content per scene, making Chunk-Based MaxInfo selection suboptimal.
-If applied globally, MaxInfo may discard crucial frames due to artificial embedding similarities across different scenes.
-To address this, we introduce Scene-Aware MaxInfo, a very simple and effective modification.
-Instead of performing complex scene detection, we uniformly split the video into 32 equal-sized chunks and apply MaxVol independently within each chunk.
-This ensures that each segment is well-represented while maintaining computational simplicity.
-Chunk-Based MaxVol Selection
-Given $n$ uniformly sampled frames, we split them into 32 contiguous chunks:
-$\mathbf{I}=\bigcup_{i=1}^{32}\mathbf{I}^{(i)},\quad\mathbf{I}^{(i)}=\{i_{j}\mid j\in\text{chunk }i\}.$ (9)
-For each chunk, we extract CLIP embeddings, apply SVD for dimension reduction, and run MaxVol to select the most representative frames.
+3.5 Chunk-Based MaxInfo
+Videos often contain multiple scenes with visually similar frames, making global frame selection suboptimal.
+Applied across the entire video, MaxInfo may also discard important frames due to spurious embedding similarities between different scenes.
+To address this, we propose Chunk-Based MaxInfo, a simple yet effective modification.
+We uniformly divide the video into $M$ equal-sized chunks and apply MaxInfo independently within each chunk.
+This ensures that every segment is adequately represented while keeping the procedure computationally efficient.
+Formally, given $n$ uniformly sampled frames, we split them into $M$ contiguous chunks:
+$\mathbf{I}=\bigcup_{i=1}^{M}\mathbf{I}^{(i)},\quad\mathbf{I}^{(i)}=\{i_{j}\mid j\in\text{chunk }i\}.$ (9)
+For each chunk, we extract CLIP embeddings, apply SVD for dimensionality reduction, and run MaxVol to select the most representative frames.
+In our experiments, we set $M=32$ for simplicity, though any choice of $M$ is possible and can be tuned to balance representation quality and computational cost.
 This approach is deliberately simple but demonstrates that even Chunk-Based scene segmentation can further enhance MaxInfo’s effectiveness.
@@
 3.6 Summary
-Our pipeline can be summarized with an Algorithm 1.
-Because our method is training-free and focuses purely on data selection, it can be plugged into diverse video-focused VLLMs to reduce cost and improve performance on long-video tasks.
+Our pipeline can be summarized with an Algorithm 1. It is a training-free and plug-and-play algorithm that can be integrated into any model of VLLMs.
 Algorithm 1 MaxInfo Block: SVD + MaxVol for Keyframe Selection
@@
 2:Embedding: Convert each frame $i_{j}$ into a [CLS] embedding:
-$q_{n}=\mathrm{flatten}\bigl{(}\mathrm{clip\_model}(i_{n})\bigr{)},\mathbf{Q}=\begin{bmatrix}q_{1}\\ q_{2}\\ \vdots\\ q_{n}\end{bmatrix}\in\mathbb{R}^{n\times d}.$
+$q_{n}=\mathrm{flatten}\bigl(\mathrm{clip}(i_{n})\bigr),\mathbf{Q}=\begin{bmatrix}q_{1}&q_{2}&\cdots&q_{n}\end{bmatrix}^{\top}$
 3:SVD Reduction: Perform truncated SVD on $\mathbf{Q}$:
 $\mathbf{Q}\approx\mathbf{U}_{r}\,\mathbf{\Sigma}_{r}\,\mathbf{V}_{r}^{\top}\quad\rightarrow\quad\mathbf{Q_{s}}=\mathbf{U}_{r}\in\mathbb{R}^{n\times r}.$
-[…]
+4:MaxVol Selection: Run $\text{rect\_maxvol}(\mathbf{Q_{s}},\mathrm{tol})$ to find pivot indices:
 $\mathrm{piv}=\text{rect\_maxvol}(\mathbf{Q_{s}},\mathrm{Tol}),$
@@
 4 Experiments
-4.1 Implementation Details
-In order to evaluate the contribution of MaxInfo to video understanding, we employed widely-used video understanding benchmarks, covering short-video tasks (such as MVBench (Li et al., 2024)) and medium-to-long video tasks (such as EgoSchema (Mangalam et al., 2023), Video-MME (Fu et al., 2024), and LongVideoBench (Shen et al., 2024)).
+Overall.
+In order to evaluate the contribution of MaxInfo to video understanding, we employed widely-used video understanding benchmarks, covering short-video tasks (such as MVBench [mvbench]) and medium-to-long video tasks (such as EgoSchema [egoschema], Video-MME [video-mme], and LongVideoBench [longvu]).
 For complete fairness, we only compared improved versions of the model against itself without MaxVol with freezing generation parameters, seed and prompt.
+4.1 Main Results
+Overall Performance.
+Table 1 present the performance gains achieved by integrating MaxInfo into existing InternVL2 [internvl], Qwen2-VL [qwen2-vl] and LLaVA-Video [llava-video] models.
+Experimental results show that MaxInfo Block exhibits significant performance improvements on a number of models.
+In particular, in the LLaVA-Video-7B and Qwen-VL-2B models, the introduction of MaxInfo Block improves the accuracy by 0.9%/1.7%, 6.4%, 3.3% and 1.4%/1.2%, 2.3%, 1.5% in VideoMME [video-mme], EgoSchema [egoschema], and LongVideoBench [longvideobench], respectively, which is significantly better than the versions without the MaxInfo block.
+This improvement not only validates the effectiveness of MaxInfo Block, but also provides new ideas for future research on video understanding tasks.
+Table 1: Comparison of VLLM with and without MaxInfo on multiple benchmarks, where wo. sub. is without subtitles and with w. sub. subtitles.
+Model Size VideoMME (wo/w-subs) Egoshcema LongVideoBench
+LLaVA-Video [llava-video] 7B 63.3/69.7(64) 57.3(64) 58.2(64)
++ MaxInfo 7B 64.2/71.464→(6,64) 63.7128→(12,64) 61.5128→(1,64)
+[…]
+LLaVA-Video [llava-video] 72B 70.5/76.9(64) 65.6(64) 61.9(64)
++ MaxInfo 72B 70.9/77.664→(6,64) 69.4128→(12,64) 64.9128→(1,64)
+[…]
+Qwen2-VL [qwen2-vl] 2B 55.6/60.4(786) 54.9(180) 47.3(256)
++ MaxInfo 2B 57.0/61.6256→(4,254) 57.2180→(12,180) 48.8256→(1,224)
+[…]
+Qwen2-VL [qwen2-vl] 7B 63.3/69.0(768) 66.7(180) 53.7(256)
++ MaxInfo 7B 62.1/70.0256→(4,254) 64.3180→(12,180) 55.7256→(1,224)
+[…]
+InternVL2 [internvl] 1B 43.0(16) 34.0(128) 43.4(128)
++ MaxInfo 1B 43.6128→(1,16) 34.1128→(1,32) 43.9128→(1,32)
+[…]
+InternVL2 [internvl] 2B 45.4(16) 46.1(128) 47.0(128)
++ MaxInfo 2B 45.0128→(1,16) 46.3128→(1,32) 47.3128→(1,32)
+[…]
+Although our results are slightly lower than the baseline on some models, we use significantly fewer frames than the baseline configuration.
+Chunck Based MaxVol.
+Table 2 evaluates MaxInfo and Chunk-Based Scene-Awareness on the MLVU benchmark, showing consistent gains over uniform sampling.
+This approach is simple and entirely training-free, highlighting the potential of even basic scene-awareness.
+These results suggest that more advanced scene segmentation techniques could yield further improvements, making scene-awareness a promising direction for video understanding.
+Table 2: Performance comparison on the MLVU benchmark between original Qwen2-VL and its variants with MaxVol and Chunk-Based. Best results in bold, second-best underlined.
+Model Size Acc.
+Qwen2-VL [qwen2-vl] 2B 52.18
++ MaxVol 2B 52.37
++ Chunk-Based 2B 52.69
+Qwen2-VL [qwen2-vl] 7B 64.32
++ MaxVol 7B 64.59
++ Chunk-Based 7B 64.82
+Comparison Baseline.
+As shown in Table 3, our proposed MaxInfo achieves strong competitiveness among training-free keyframe extraction methods and further demonstrates comparable or superior performance relative to existing state-of-the-art approaches.
+Here, * indicates that our method does not require a predefined number of frames, but dynamically selects key frames based on the model’s initial inference frame count and the video’s information content.
+For example, on Video-MME and LongVideoBench, Qwen-MaxInfo processes an average of 180 and 170 frames, respectively, compared to 768 and 256 frames in the base Qwen model.
+Similarly, LLaVA-Video-MaxInfo uses 64 frames on both datasets, while the base LLaVA-Video model uses 58 and 51 frames, respectively.
+Table 3: This table compares the performance of VLM with current state-of-the-art (SOTA) open-source and proprietary models, as well as key-frame selection methods, on video benchmark key-frame selection tasks.
+First place: bold, second place: underline, third place: italic.
+Model LLM size #Frames VideoMME (wo sub.) EgoSchema LongVideoBench (val.)
+Short Medium Long Overall
+1.3min 9min 41min 17min 3min 12min
+Proprietary Models
+GPT4-o [openai2024gpt4o] - 1fps 77.1 62.1 59.2 66.2 - 66.7
+Gemini-1.5-Pro [team2024gemini] - 1fps 82.3 75.3 67.5 75.7 - 64.0
+Open Source Models
+VideoChat2 [mvbench] 7B 16 48.3 37.0 33.2 39.5 - -
+ShareGPT4Video [chen2024sharegpt4video] 8B 16 - - 37.9 43.6 - -
+VideoLLaMA2 [cheng2024videollama] 7B 32 56.0 45.4 42.1 47.9 - -
+LongVILA [xue2024longvila] 8B 128 60.2 48.2 38.8 49.2 - -
+8B 256 61.8 49.7 39.7 50.5 - -
+Qwen2-VL [qwen2] 7B 8 65.0 50.7 45.3 53.7 53.5 -
+Key-frames Selection Methods
+LongVU [longvu] 7B 1fps 64.7 58.2 59.5 60.6 67.6 -
+LongVA [zhang2024long] 7B 16 59.0 46.6 43.6 49.7 - -
+64 61.4 50.9 45.0 52.4 - -
+128 61.1 50.4 46.2 52.6 - -
+384 60.3 48.9 46.1 51.8 - -
+Chat-Univi-v1.5 [Chat-univi] 7B 64 51.2 44.6 41.8 45.9 - -
+AKS (LLaVA-Video) [tang2025adaptive] 7B 64 - - - 65.3 - 62.7
+M-LLM (Qwen2-VL) [hu2025m] 7B - 69.6 54.1 51.9 58.7 65.9 -
+BOLT (LLaVA-OneVision) [liu2025bolt] 7B 32 70.1 60.0 49.6 59.9 64.0 59.6
+AdaReTaKe (Qwen2-VL) [wang2025adaretake] 7B - - - 56.4 64.2 - 57.2
+AdaReTaKe (LLaVA-Video) [wang2025adaretake] 7B - - - 53.9 64.0 - 59.6
+GenS (Qwen2-VL [yao2025generative]) 7B 50 - - - - - 58.7
+ViLaMP [cheng2025scaling] 7B 1fps - - 58.4 67.7 - 60.2
+Frame-Voyager [yu2024frame] 8B 128 (16) 67.3 56.3 48.9 57.5 - -
+VideoTree [videotree] GPT-4 avg. 62 - - 54.2 - 61.1 -
+Q-Frame (GPT-4o) [zhang2025q] GPT-4o 8 63.8 69.9 63.8 57.6 - 58.6
+VSLS (GPT-4o) [guo2025logic] GPT-4o 32 71.9 61.9 55.2 63.0 - 63.4
+VSLS (InternVL2.5-78B) [guo2025logic] 78B 8 59.0 57.5 57.7 58.1 - 64.5
+MaxInfo (ours)
+MaxInfo (InternVL2) 1B * - - - - - 43.9
+MaxInfo (Qwen2-VL) 7B * 72.5 62.0 51.8 62.1 64.3 55.7
+MaxInfo (LLaVA-Video) 7B * 74.6 63.3 54.6 64.2 63.7 61.5
+MaxInfo (LLaVA-Video) 72B * 80.2 67.7 62.7 70.2 69.4 64.9
+4.2 Fast vs. Slow Version Comparisons
+As shown in Table 4, the slow version outperforms the fast version in most cases, especially when processing long videos.
+Its initial oversampling mechanism provides MaxInfo with a richer selection space, which significantly improves performance.
+However, experiments have also shown that the fast version may outperform the slow version in certain benchmarks or when the number of initial samples is small.
+The MaxInfo block time is very short, so latency is almost negligible and more details on latency and GPUs memory and time.
+Table 4: Performance of Fast vs.
+Slow MaxInfo variants on LLaVA-Video-7B (LongVideoBench).
+I/O: input/output frames.
+All results are reproduced, except those marked with *, which are copied from the corresponding papers [llava-video].
+Model I max. frames O frame range Avg. frames Encoding + MaxInfo Time Acc.
+InternVL2 1B [internvl] 32 - 32 - 43.38
++ MaxInfofast 32 (1, 32) 30 0.179 + 0.0126 s 43.60 +0.22%
++ MaxInfoslow 64 (1, 16) 16 0.339 + 0.0215 s 44.65 +1.05%
+InternVL2 2B [internvl] 32 - 32 - 46.97
++ MaxInfofast 32 (1, 32) 30 0.179 + 0.0126 s 47.27 +0.30%
++ MaxInfoslow 64 (1, 16) 16 0.339 + 0.0215 s 46.82 -0.15%
+LLaVA-Video 7B [llava-video] 64 - 64 - 58.2
++ MaxInfofast 64 (1, 64) 52 0.339 + 0.0215 s 60.21 +2.01%
++ MaxInfoslow 128 (1, 64) 58 0.624 + 0.0421 s 61.48 +3.28%
+LLaVA-Video 72B [llava-video] 64 - 64 - 61.9
++ MaxInfofast 64 (1, 64) 52 0.339 + 0.0215 s 65.37 +3.47%
++ MaxInfoslow 128 (1, 64) 58 0.624 + 0.0421 s 64.92 +3.02%
+4.3 Ablation Study
+To further evaluate the effectiveness of MaxInfo, we conducted a series of ablation experiments focusing on two key aspects: the impact of the choice of visual encoder and the influence of key hyperparameters in the MaxInfo module, particularly tolerance and rank.
+4.3.1 Vision Encoder Impact
+As shown in Table 5, we evaluated CLIP, DINOv2, and SigLIP as visual encoders.
+DINOv2 performed comparably to CLIP, despite lacking vision-language alignment.
+However, SigLIP outperformed both CLIP and DINOv2, likely due to its stronger language-vision connectivity.
+Interestingly, the larger SigLIP model underperformed, suggesting that more complex encoders require careful hyperparameter tuning (e.g., tolerance and rank) for optimal frame selection.
+Table 5: Visual encoder ablation for MaxInfo. Best in bold, second-best underlined. Based on LLaVA-Video-7B [llava-video].
+Model Visual Encoder Param. Acc.
+LLaVA-Video [llava-video] CLIP-ViT-Large 427.9M 58.94
+LLaVA-Video [llava-video] CLIP-ViT-Base 149.6M 58.79
+LLaVA-Video [llava-video] DINO2-base 86.6M 58.94
+LLaVA-Video [llava-video] DINO2-large 304.4M 58.86
+LLaVA-Video [llava-video] SigLIP-base-224 203.2M 59.76
+LLaVA-Video [llava-video] SigLIP-base-384 878.0M 59.24
+4.3.2 Impact of MaxInfo Block Hyperparameters
+This section examines the effect of MaxInfo parameters (rank R and tolerance Tol) as well as the initial number of samples on the final model accuracy.
+Effect of Different Tolerances and Ranks on the Model with Fixed Sampling.
+We conducted a series of tests with fixed benchmarks, model settings, and an initial pool of frames (e.g., $n^{*}=96$).
+As shown in Figure 3, the best observed result achieved a 3.3% improvement over the base LLaVA-Video-7B model with $\text{Tol}=0.3$ and $\text{R}=8$.
+From these experiments, we derived the following guidelines:
+  1. 1.
+    Performance Sensitivity: Our experiments show that performance is most sensitive to Tol values in the range $\text{Tol}\in[0.15,0.60]$.
+Beyond this range, improvements plateau or regress due to over-pruning or under-pruning.
+  2. 2.
+    Convergence: The model will converge when setting $\text{R}\in[12,15],\text{Tol}\in[0.3,0.45]$, this means that our choice of hyperparameters is not intractable.
+Figure 3: Effect of initial sampling on MaxInfo performance for LlaVa-Video 7B model.
+Effect of Sampling on Accuracy We analyzed how varying the initial number of sampled frames impacts accuracy, while keeping the hyperparameters fixed (R=8, Tol=0.15).
+As shown in Figure 4, accuracy initially improves with the addition of more frames, but beyond a certain threshold, it begins to plateau or slightly decline.
+Our key observations are as follows:
+  1. 1.
+    Increasing the initial number of frames provides more diverse information, enabling MaxVol to select better keyframes, but the information converges.
+  2. 2.
+    There exists an optimal trade-off between the initial frame count and computational cost. In our tests, 128 as initial frames yielded the best accuracy for the LLaVA-Video-7B model.
+Figure 4: Effect of Initial Sampling on MaxInfo. Starting from $n^{*}$ sampled frames, the MaxInfo Block selects up to 64 informative frames for further processing.
+4.3.3 Case Study
+Figure 5 presents a long-video example from the Video-MME [video-mme] dataset, qualitatively illustrating the effectiveness of the proposed MaxInfo Block.
+We compare frame selection by MaxInfo Block with Uniform Sampling and observe that the keyframes chosen by MaxInfo are more closely aligned with the manually annotated Ground Truth–related frames.
+Figure 5: Qualitative: (a) MaxInfo vs Uniform Sampling with GT-aligned frames; (b) CLIP scores show MaxInfo’s answer coverage in single samples.
+5 Conclusion
+In this work, we introduced MaxInfo – a training-free method for selecting the most informative frames from videos, improving VLLMs inference.
+Our results consistently demonstrate that informative frame selection outperforms uniform sampling, leading to improved performance of state of the art VLLMs (LLaVA-Video, InternVL and Qwen2-VL with different sizes) across multiple benchmarks.
+For example, MaxInfo achieves a 3.28% improvement on LongVideoBench and a 6.4% improvement on EgoSchema for LLaVA-Video-7B.
+The Slow Version of MaxInfo improves LLaVA-Video-72B performance by 3.47% on LongVideoBench.
+Beyond demonstrating empirical gains, we believe our work will encourage the community to focus more on frame selection strategies, an often-overlooked aspect of video understanding.
+Additionally, we have shown that even minimal refinements, such as chunk-wise MaxVol in our Scene-Aware MaxInfo, can further enhance results, demonstrating that simple adjustments can lead to meaningful improvements.
+Finally, we hypothesize that training VLLMs with informative frame sampling, rather than simple uniform frame selection, could further enhance their capabilities when later used with inference-time MaxInfo techniques.
+We hope this work serves as a foundation for future research into more efficient, information-aware video sampling strategies for large-scale multimodal learning.
+6 Acknowledgments
+Innopolis University authors were supported by the Research Center of the Artificial Intelligence Institute at Innopolis University.
+Financial support was provided by the Ministry of Economic Development of the Russian Federation (No.
+25-139-66879-1-0003).
+Appendix A Implementation Details
 We mostly focus on longer videos because better frame selection plays a bigger role in longer, more complex videos, whereas shorter ones intuitively work well with uniform sampling due to their lower information content and complexity.
@@
 For the general multiple-choice question-answering evaluation, we follow the official guidelines to construct the instructions using the provided questions and options.
-We added a prompt to the question and options like ”Respond with only the letter (A, B, C, or D) of the correct option.” for LongVideoBench (Wu et al., 2024), Video-MME (Fu et al., 2024), MLVU (Zhou et al., 2024) and MVBench (Li et al., 2024) or ”Answer with the option’s letter from the given choices directly and only give the best option.” for EgoSchema (Mangalam et al., 2023).
-We follow the original benchmarks setup to calculate the final scores, and we also align our evaluation protocols with other evaluation toolkits, such as lmms-eval (Zhang et al., 2024a).
+We added a prompt to the question and options like ”Respond with only the letter (A, B, C, or D) of the correct option.” for LongVideoBench [longvideobench], Video-MME [video-mme], MLVU [mlvu] and MVBench [mvbench] or ”Answer with the option’s letter from the given choices directly and only give the best option.” for EgoSchema [egoschema].
+We follow the original benchmarks setup to calculate the final scores, and we also align our evaluation protocols with other evaluation toolkits, such as lmms-eval [zhang2024lmmseval].
 To ensure the reproducibility of our results, we have included the main hyperparameters used for all benchmarks and estimated models in the results tables, such as tolerance and rank for the MaxInfo algorithm, the number of sampled frames, and the number of initial frames (before MaxInfo).
-4.2 Main Results
-Overall Performance.
-Table 1, Table 2 and Table 3 present the performance gains achieved by integrating MaxInfo into existing InternVL (Chen et al., 2024), Qwen2-VL (Wang et al., 2024a) and LlaVA-Video (Zhang et al., 2024c) models.
-The experimental results show that MaxInfo Block exhibits significant performance improvement on a wide range of models.
-In particular, in the LLaVA-Video-7B and Qwen-VL-2B models, the introduction of MaxInfo Block improves the accuracy by 0.9%/1.7%, 6.4%, 3.3% and 1.4%/1.2%, 2.3%, 1.5% in VideoMME, EgoSchema, and LongVideoBench, respectively, which is significantly better than the versions without the block.
-This improvement not only validates the effectiveness of MaxInfo Block, but also provides new ideas for future research on video understanding tasks.
-Table 1: Comparing Video-LLMs with and without MaxInfo as an additional module.
-Except for ours (-MaxInfo) and the models with *, all results are copied from the related papers of benchmarks or models.
-Base models uses uniform frames sampling.
-For the Video-MME benchmark, we report results under two standard settings: without subtitles (wo.
-sub.) and with subtitles (w.
-sub.).
-The % accuracy sign is omitted for clarity.
-Model Size VideoMME (wo/w-subs) Egoshcema LongVideoBench
-LLaVA-Video (Zhang et al., 2024c) 7B 63.3/69.7(64) 57.3(64) 58.2(64)
-+ MaxInfo 7B 64.2/71.464→(6,64) 63.7128→(64,64) 61.5128→(1,64)
-$\triangle$ +0.9%/+1.7% +6.4% +3.3%
-LLaVA- (Zhang et al., 2024c) 72B 70.5/76.9(64) 65.6(64) 61.9(64)
-+ MaxInfo 72B 70.2/77.664→(6,64) 69.4128→(64,64) 64.9128→(64,1)
-$\triangle$ -0.3%/+0.7% +3.8% +3%
-Qwen2-VL (Wang et al., 2024a) 2B 55.6/60.4(786) 54.9(180) 47.3(256)
-+ MaxInfo 2B 57.0/61.6256→(254,4) 57.2180→(180,12) 48.8256→(224,1)
-$\triangle$ +1.4%/+1.2% +2.3% +1.5%
-Qwen2-VL (Wang et al., 2024a) 7B 63.3/69.0(768) 66.7(180) 53.7(256)
-+ MaxInfo 7B 62.1/70.0256→(254,4) 64.3180→(180,12) 55.7256→(224,1)
-$\triangle$ -1.2/+1.0% +2.4% +2.0%
-Table 2: Comparison of the InternVL2 model series with and without MaxInfo as an additional module.
-InternVL2 uses uniform frame sampling and has a limited context size.
-Therefore, the maximum number of frames for evaluation across all benchmarks was limited to 32.
-The maximum number of input and output frames for the MaxInfo module is set to 128 and 32 respectively.
-For the LongVideoBench benchmark, we report results under a default setting for video LLMs, i.e.
-without subtitles.
-The % accuracy sign is omitted for clarity.
-Model Size MVBench Egoschema LongVideoBench
-InternVL2 (Chen et al., 2024) 1B 55.42 34.02 43.38
-+ MaxInfo(128,32) 1B 56.07 34.11 42.78
-$\triangle$ +0.65% +0.09% -0.6%
-InternVL2 (Chen et al., 2024) 2B 59.72 46.10 43.97
-+ MaxInfo(128,32) 2B 59.02 46.33 47.27
-$\triangle$ -0.7% +0.23% +3.3%
-Table 3: Performance comparison between MaxInfo Block and other key frame selection methods on the Video-MME-Long benchmark.
-Results are measured without adding subtitles.
-$Q$ refers to the Qwen2-VL-7B model and $L$ refers to the LLaVA-Video-7B model.
-The % accuracy sign is omitted for clarity.
-Model Size frames Acc.
-Video-Tree (Wang et al., 2024c) - - 54.2/-
-Long-VU (Shen et al., 2024) 7B 1fps -/60.6
-Chat-Univi (Jin et al., 2024a) 7B 64 35.8/-
-Q+MaxInfo 7B 256 $\rightarrow$ (254, 4) 51.8/63.9
-L+MaxInfo 7B 64 $\rightarrow$ (64, 6) 54.6/65.9
-MLVU Benchmark.
-Table 4 evaluates both MaxInfo and the benefits of Chunk-Based Scene-Awareness on the MLVU benchmark, demonstrating consistent improvements over uniform sampling.
-While these improvements are modest, it is important to highlight that the method is exceptionally simple and entirely training-free.
-This simplicity underscores the potential of even basic scene-awareness to enhance performance, suggesting that more sophisticated scene detection techniques could lead to further gains.
-Such results indicate that scene-awareness is a promising direction for improving video understanding tasks.
-Table 4: Comparison of performance on the MLVU benchmark between the original versions of the models in the Qwen-VL family and their variants with the MaxInfo and Naive Scene-Aware (NSA) additional modules.
-The original models use uniform sampling with a maximum of 128 frames.
-The additional modules use from 128 to 1024 frames per input.
-The best results are highlighed in bold, second best are underlined.
-Model Size Acc.
-Qwen2-VL (Wang et al., 2024a) 2B 52.18
-+ MaxVol 2B 52.37
-+ Chunk-Based 2B 52.69
-Qwen2-VL (Wang et al., 2024a) 7B 64.32
-+ MaxVol 7B 64.59
-+ Chunk-Based 7B 64.82
-4.3 Performance Analysis: MaxInfo vs. Uniform Sampling
+Appendix B Additional Experiments and Details
+To further assess the impact of MaxInfo, we evaluate its performance with an additional set of models [wang2025internvl3], [yao2024minicpmv] on the LongVideoBench and Video-MME benchmarks.
+B.1 Applying MaxInfo to recent models
+The results in Table 6 show that MaxInfo consistently improves model performance across both benchmarks, suggesting that precise frame selection is particularly important for long-video tasks.
+Table 6: Adaptation of MaxInfo to current new long video understanding models
+Model Size Frame Interval Avg. frames LongVideoBench.
+MiniCPM [yao2024minicpmv] 9B 128 128 56.17
++ MaxInfo 9B [8, 82] 56 59.61
+[…]
+InternVL3.5 [wang2025internvl3] 1B 16 16 47.7
++ MaxInfo 1B [1, 16] 16 49.0
+[…]
+InternVL3.5 [wang2025internvl3] 8B 16 16 57.4
++ MaxInfo 8B [1, 16] 16 59.0
+[…]
+InternVL3.5 [wang2025internvl3] 38B 16 16 60
++ MaxInfo 38B [1, 16] 16 61.6
+[…]
+B.2 Performance Analysis: MaxInfo vs. Uniform Sampling
 To better understand the strengths and trade-offs of MaxInfo, we analyzed per-task accuracy across multiple benchmarks.
-Figure 3: Accuracy comparison between Uniform Sampling and MaxInfo on Video-MME (Qwen2-VL-2B).
-Figure 4: Accuracy comparison between Uniform Sampling and MaxInfo on LongVideoBench (Qwen2-VL-2B).
-Figure 5: Accuracy comparison between Uniform Sampling and MaxInfo on MLVU (Qwen2-VL-7B) with 64 final and 1024 initial frames.
-Our results, as shown in Figure 3, Figure 4, and Figure 5, indicate that MaxInfo outperforms uniform sampling on tasks requiring high information density (e.g., counting, summarization, and spatial reasoning).
-However, for tasks requiring fine-grained temporal reasoning (e.g., object recognition and action perception), uniform sampling retains a slight advantage.
-These findings highlight a fundamental trade-off between information maximization and temporal continuity, suggesting potential directions for hybrid selection methods.
-4.4 Fast vs. Slow Version Comparisons
-As Table 5 shows, the slow version outperforms the fast version in most cases, particularly when processing longer videos.
-The initial oversampling mechanism in the slow version provides MaxInfo with a richer selection space, which significantly improves performance.
-However, experimental results indicate that the fast version may outperform the slow version in specific cases, depending on the benchmark type and the number of initial samples.
-Nevertheless, in scenarios with limited computational resources or strict latency requirements, the fast version remains a practical and reliable choice due to its efficiency.
-This flexibility allows the two versions to adapt to different task requirements, offering users more options.
-Table 5: Comparison of the impact of the slow and fast versions of the MaxInfo module on the performance of the LLaVA-Video-7B model on the LongVideoBench.
-All results are reproduced, except those marked with *, which are copied from the corresponding papers (Zhang et al., 2024c).
-Model Input max. frames Output frames range Avg. frames Acc.
-InternVL2 1B (Chen et al., 2024) 32 - 32 43.38
-+ MaxInfofast 32 (1, 32) 30 43.6 +0.22%
-+ MaxInfoslow 64 (1, 16) 16 44.65 +1.05%
-InternVL2 2B (Chen et al., 2024) 32 - 32 46.97
-+ MaxInfofast 32 (1, 32) 30 47.27 +0.73%
-+ MaxInfoslow 64 (1, 16) 16 46.82 -0.15%
-LLaVA-Video 7B (Zhang et al., 2024c) 64 - 64 58.2
-+ MaxInfofast 64 (1, 64) 52 60.21 +2.01%
-+ MaxInfoslow 128 (1, 64) 58 61.48 +3.28%
-LLaVA-Video 72B (Zhang et al., 2024c) 64 - 64 61.9
-+ MaxInfofast 64 (1, 64) 52 65.37 +3.47%
-+ MaxInfoslow 128 (1, 64) 58 64.92 +3.02%
-4.5 Ablation Study
-To further evaluate the effectiveness of MaxInfo, we conducted a series of ablation experiments focusing on two key aspects: the impact of the choice of visual encoder and the influence of key hyperparameters in the MaxInfo module, particularly tolerance and rank.
-4.5.1 Vision Encoder Impact
-Table 6 presents an ablation study on different visual encoders used for frame representation in MaxInfo.
-Table 6: An ablation study on the choice of a visual encoder to obtain the visual representation of the frames in the MaxInfo approach.
-The best results are highlighed in bold, second best are underlined.
-Based on LLaVA-Video-7B (Zhang et al., 2024c)
-Model Size Visual encoder Param. Acc.
-LLava-Video 7B CLIP-ViT-Large 427.9 M. 58.94
-LLava-Video 7B CLIP-ViT-Base 149.6 M. 58.79
-LLava-Video 7B DINO2-base 86.6 M. 58.94
-LLava-Video 7B DINO2-large 304.4 M. 58.86
-LLava-Video 7B SigLIP-base-224 203.2 M. 59.76
-LLava-Video 7B SigLIP-base-384 878 M. 59.24
-We evaluated CLIP, DINOv2, and SigLIP as visual encoders.
-DINOv2 performed comparably to CLIP, despite lacking vision-language alignment.
-However, SigLIP outperformed both CLIP and DINOv2, likely due to its stronger language-vision connectivity.
-Interestingly, the larger SigLIP model underperformed, suggesting that more complex encoders require careful hyperparameter tuning (e.g., tolerance and rank) for optimal frame selection.
-4.5.2 Impact of MaxInfo Block Hyperparameters
-This section examines the effect of MaxInfo parameters (rank R and tolerance Tol) as well as the initial number of samples on the final model accuracy.
-Effect of Different Tolerances and Ranks on the Model with Fixed Sampling.
-We conducted a series of tests with fixed benchmarks, model settings, and an initial pool of frames (e.g., $n^{*}=96$).
-As shown in Figure 6, the best observed result achieved a 3.3% improvement over the base LLaVA-Video-7B model with $\text{Tol}=0.3$ and $\text{R}=8$.
-Figure 6: Effect of initial sampling on MaxInfo performance for LlaVa-Video 7B model (fixed R=8, Tol=0.15).
-From these experiments, we derived the following guidelines:
-  1. 1.
-    Large Rank, Moderate Tol: When the Rank $R$ is large, a moderate Tolerance (Tol) helps avoid selecting overly similar frames, optimizing the trade-off between diversity and redundancy.
-  2. 2.
-    Balanced Selection: An excessively large $R$ combined with a very small Tol may cause the frame set to grow excessively, potentially exceeding the model’s capacity.
-Thus, $R$ and Tol should be selected based on the maximum feasible input frame count for a given LLM.
-  3. 3.
-    Performance Sensitivity: Our experiments show that performance is most sensitive to Tol values in the range $\text{Tol}\in[0.1,0.6]$.
-Beyond this range, improvements plateau or regress due to over-pruning or under-pruning.
-  4. 4.
-    Convergence: When setting $R=12$ and varying Tol values to 0.30, 0.45, and 0.60, the best result converged at 61.1%.
-Effect of Sampling on Accuracy (Fixed Rank and Tolerance).
-We analyzed how varying the initial number of sampled frames impacts accuracy, while keeping the hyperparameters fixed (R=8, Tol=0.15).
-As shown in Figure 7, accuracy initially improves with the addition of more frames, but beyond a certain threshold, it begins to plateau or slightly decline.
-Our key observations are as follows:
-Figure 7: Comparing the effect of initial sampling on the MaxInfo Block, n represents the initial number of frames sampled, and after passing through the MaxInfo Block, the number of frames is increased to a maximum of 64.
-  1. 1.
-    Increasing the initial number of frames provides more diverse information, enabling MaxVol to select better keyframes.
-  2. 2.
-    Beyond a certain number of frames, accuracy slightly declines due to excessive redundancy, which hampers MaxVol’s efficiency in keyframe selection.
-  3. 3.
-    There exists an optimal trade-off between the initial frame count and computational cost. In our tests, 128 as initial frames yielded the best accuracy for the LLaVA-Video-7B model.
-5 Conclusion
-In this work, we introduced MaxInfo – a training-free method for selecting the most informative frames from videos, improving VLLMs inference.
-Our results consistently demonstrate that informative frame selection outperforms uniform sampling, leading to improved performance of state of the art VLLMs (LLaVA-Video, InternVL and Qwen2-VL with different sizes) across multiple benchmarks.
-However, the improvements are relatively small, which is expected given the training-free nature and the simplicity of our approach.
-Beyond demonstrating empirical gains, we believe our work will encourage the community to focus more on frame selection strategies, an often-overlooked aspect of video understanding.
-Additionally, we have shown that even minimal refinements, such as chunk-wise MaxVol in our Scene-Aware MaxInfo, can further enhance results, demonstrating that simple adjustments can lead to meaningful improvements.
-Finally, we hypothesize that training VLLMs with informative frame sampling, rather than simple uniform frame selection, could further enhance their capabilities when later used with inference-time MaxInfo techniques.
-We hope this work serves as a foundation for future research into more efficient, information-aware video sampling strategies for large-scale multimodal learning.
-Impact Statement
+Our results, as shown in Figure 6, indicate that MaxInfo performs superiorly in high information density tasks such as counting, summarizing and spatial reasoning, while uniform sampling has a slight advantage in tasks that rely on temporal continuity, reflecting the key trade-off between information maximization and temporal consistency.
+Figure 6: Accuracy comparison between Uniform Sampling and MaxInfo across three benchmarks.
+B.3 Comparison with CLIP baseline
+As shown in Table 7, we compare the experimental results of two keyframe extraction strategies based on the QwenVL2-2B model on the LongVideoBench benchmark: the CLIP-Based thresholding method and the MaxInfo module method.
+Both methods extract the same number of frames in the initial phase, so the encoding time is kept the same, where the similarity threshold of the CLIP-Based method is set to 0.5.
+The results show that the MaxInfo module outperforms the CLIP-Based method in terms of the overall performance in keyframe selection.
+Table 7: Performance comparison on LVBench.
+Model Method Accuracy
+QwenVL2-2B CLIP-Based 44.3
+QwenVL2-2B CLIP-Based + MaxInfo 44.5
+QwenVL2-2B MaxInfo + CLIP-Based 43.8
+QwenVL2-2B MaxInfo 48.8
+In addition, we also explored combining the CLIP-Based method with MaxInfo module.
+The experiments show that MaxInfo is able to improve the overall information quality of the input sequences, and its information maximization strategy plays a key role in frame selection, which further enhances the performance of the model.
+CLIP-Based loses a lot of semantic information, which can lead to performance degradation of the model.
+In order to further evaluate whether MaxInfo will lose the key frames related to the problem, we compare MaxInfo with the Uniform Sampling method under the CLIP Score metric.
+The experimental results shown in Table 8 that MaxInfo does not miss the frames related to the semantics of the problem, and is able to retain the semantic relevance effectively.
+Table 8: CLIP score comparison between uniform and MaxInfo sampling.
+Sampling Method CLIP-score
+Uniform 0.37
+MaxInfo 0.39
+B.4 Qualitative comparison with uniform sampling
+We randomly selected 50 video samples in LongVideoBench and calculated the cosine similarity between the frames selected by MaxInfo and the cosine similarity between the frames obtained by uniform sampling.
+Figure 7 shows the distribution of cosine similarity for the same number of frames.
+It is clear that MaxInfo produces a more diverse distribution like a low similarity offset compared to uniform sampling, highlighting its ability to capture more diverse visual content.
+Figure 7: Similarity distribution between neighbouring frames ($frame_{i}$ and $frame_{i+1}$).
+As shown in Figure 8, we plotted 200 sampled data points to improve visual clarity. The results show that our MaxInfo module exhibits higher diversity in frame selection compared to uniform sampling.
+Figure 8: CLIP similarity between neighboring frames selected by MaxInfo module.
+B.5 Computational Efficiency: Time and Memory Consumption
+When processing long videos, the LLM is the most resource-intensive component of VLLMs due to its parameter count and the quadratic complexity of attention with respect to input length.
+Since most of the context is occupied by visual tokens from frames, our MaxInfo method reduces this load by selecting keyframes.
+Importantly, MaxInfo requires minimal and constant memory and preprocessing time, independent of LLM size, and remains significantly lighter than uniformly sampling all frames.
+Time Complexity.
+To evaluate the latency overhead of MaxInfo in practice, we measured its runtime with the Qwen2-VL model.
+As shown in Table 9, the runtime of MaxInfo is almost negligible compared to the inference time of the VLLM itself, confirming that MaxInfo is a lightweight and efficient frame selection mechanism.
+The initial VLLM time means the inference time of the 512 frames of information directly into the Qwen2-VL model.
+The frame count selected by MaxInfo Block is adaptive to the information content of the input.
+For near-static videos (low information density), MaxInfo drastically reduces the number of processed frames.
+Consequently, VLLMs + MaxInfo Block may achieve lower time compared to the initial VLLMs + MaxInfo Block configuration.
+The times reported in the table represent an upper bound; in practice, the reduced number of frames can lead to several-fold speedups on certain tasks.
+All experiments were conducted on an A100 GPU.
+Table 9: Runtime of different pipeline components, based on Qwen2-VL. Frame size = 512 (UP is Upper Bound).
+Model Size CLIP (s) MaxVol (s) VLLMs (s) VLLMs + MaxInfo (UP)
+2B 0.296 0.0109 2.979 $\leq$ 3.285
+7B 0.296 0.0109 5.372 $\leq$ 5.679
+72B 0.296 0.0109 30.737 $\leq$ 31.044
+We also analyzed the running time of the MaxVol algorithm alone, including its chunk-based variant, under different initial numbers of frames, as shown in Table 10.
+The experimental results show that the running time of MaxVol remains low across settings, with minimal impact on the overall inference efficiency.
+Table 10: MaxVol algorithm runtime (excluding image encoding time) for different input sizes.
+Method Input Size MaxVol Time (s)
+MaxInfo 128 0.0044
+MaxInfo 256 0.0053
+MaxInfo 512 0.0109
+Chunks-Based MaxInfo $32\times 32$ 0.0375
+Then we estimated CUDA inference time across different VLLM sizes which is shown in Figure 9.
+The overhead of MaxInfo remains small and nearly constant, while the overall inference time grows with model size, demonstrating that MaxInfo adds minimal cost compared to the savings from reduced visual tokens.
+For small models (up to 8B parameters), the relative benefit is limited since inference cost is low.
+However, for larger models (26B–76B), MaxInfo provides clear efficiency gains by substantially reducing the number of visual tokens, making its impact especially pronounced for long-video tasks where input length dominates computational cost.
+Figure 9: CUDA inference time across different VLLM sizes. The preprocessing cost of MaxInfo remains small and nearly constant, while overall inference time increases with model size.
+Figure 10: The comparison for memory performance on the GPU for InternVL2 models with and without MaxInfo module. The dashed line shows the CUDA memory requirements for MaxInfo.
+Memory Consumption.
+Secondly, we precisely evaluated memory consumption of out approach.
+As shown in Figure 10, MaxInfo’s CUDA memory usage remains constant for a fixed number of initial frames and grows much more slowly than uniform sampling as LLM size increases.
+In summary, our analysis of time and memory efficiency shows that MaxInfo introduces only negligible overhead while substantially reducing the computational burden of processing long videos.
+Its constant preprocessing cost and slower growth in memory usage make MaxInfo particularly advantageous for large-scale VLLMs, with the benefits becoming most pronounced for models exceeding 10B parameters.
+Appendix C Theoretical Justification
+Definition 1.
+Definition of the maximum volume of the video frame feature matrix.
+We consider a matrix $Q\in\mathbb{C}^{N\times r}$, where each row represents the CLIP or SigLIP etc.
+feature of a video frame, ordered sequentially in time, where N denotes the number of frames and r denotes the dimension of the feature.
+We aim to identify a submatrix $\hat{Q}\in\mathbb{C}^{K\times r}$ of the original matrix $S$, such that $\hat{S}$ closely approximates $S$ in terms of matrix volume, thereby preserving its essential structural information.
+To obtain the submatrix $\hat{Q}$, we introduce a coefficient matrix $C$ based on the minimum-norm linear combination as Equation 10
+$\tilde{C}\hat{Q}=\tilde{Q}$ (10)
+Here, $\tilde{Q}$ denotes a set of sample rows selected from the original matrix $Q$ for reconstruction.
+By solving for $\tilde{C}$, we can approximate the reconstruction of $\tilde{Q}$ using only the representative rows in $\hat{Q}$.
+In addition, it is shown that the selected K rows are the most representative of the video frame information.
+Solving. The submatrix $\hat{Q}\in\mathbb{C}^{K\times r}$ provides an approximation of the original matrix $Q\in\mathbb{C}^{N\times r}$ within a tolerance $\tau$.
+We start with an initial submatrix $\hat{Q}\in\mathbb{C}^{M\times r}$ and add a row $Q_{i}\in\mathbb{C}^{1\times r}$ to each iteration to bring the expanded submatrix up to speed in the sense of volume.
+The updating process can be expressed as follows Equation 11 and the volume of the updated matrix can be defined as Equation 12
+$\hat{Q}\leftarrow\begin{bmatrix}\hat{Q}\\ Q_{i}\end{bmatrix}$ (11)
+$\text{Vol}(\hat{Q})_{\text{new}}=\text{Vol}(\hat{Q})_{\text{old}}\cdot\sqrt{1+\|\tilde{C}_{i}\|_{2}^{2}}$ (12)
+where $Q_{i}$ is the row selected from the original matrix $Q\in\mathbb{C}^{N\times r}$ that currently boosts the volume of the submatrix the most.
+Repeat this process iteratively until the conditional Equation 13 is satisfied or the target number of $K$ rows is reached.
+$\|\tilde{C}_{i}\|_{2}\leq\tau$ (13)
+Proof of maximum information entropy.
+To justify our approach, we use differential entropy as an information measure.
+Suppose our normalized frame embeddings form a matrix $S$.
+The differential entropy of a uniform distribution over the convex hull $\mathcal{C}(S)$ is given by the following Equation 14.
+$H_{\max}(S)=\ln(\text{Vol}(\mathcal{C}(S)))$ (14)
+where $\text{Vol}(\mathcal{C}(S))$ is the volume of the convex hull formed by selected embeddings. Classical results show the following Equation 15.
+$\text{Vol}(\mathcal{C}(S))=\kappa\sqrt{\det(S^{\top}S)}$ (15)
+for some constant $\kappa>0$. Thus we get Equation 16
+$H_{\max}(S)=\ln V(S)+\text{constant}$ (16)
+where $V(S)=\sqrt{\det(S^{\top}S)}$. Since MaxVol maximizes $V(S)$, it maximizes the upper bound on differential entropy, ensuring that selected frames are more informative.
+In summary, we can theoretically select the most representative frame information.
+The feature matrices corresponding to the selected frames have good linear independence under the constraint of the tolerance parameter $\tau$, thus constituting an approximately optimal subset of the representation.
+This process achieves our goal of information maximization, i.e.
+preserving the most critical structural information while compressing redundancy.
+Appendix D Societal Impacts
 This work introduces a training-free framework for improving frame sampling in Vision-Language Large Models (VLLMs), enhancing video understanding tasks.
-Such advancements are crucial for applications in education, accessibility, and public safety, where accurate and efficient video analysis can have a meaningful impact.
-However, improved video analysis capabilities may also raise ethical concerns, including potential misuse in surveillance, privacy violations, or biases in automated systems.
-Ensuring responsible deployment with a focus on fairness and transparency is essential to mitigate these risks.
-In summary, while our approach offers significant benefits, its adoption must adhere to ethical principles to promote equitable and responsible use.
-Appendix A Appendix
-A.1 Visualisation of MaxInfo Block results
-Figure 8: MaxInfo Block sampled frames.
-As shown Figure 8, MaxInfo Block can help us to reduce the redundancy of video information, better understand the video content and improve the model performance by selecting the most informative and representative key frames among multiple frames.
-A.2 Additional comparisons with other models
-We compared the performance of VLLMs of the same 7B size, as shown in Table 7, such as Open Source Model, Proprietary Models and Train-Free Model or Key Frames selection, on Video-MME Benchmark.
-Table 7: Comparison of model results on Video-MME Benchmark.
-Model LLM Size Frames Overall (WO/W-Subs) Short (WO/W-Subs) Medium (WO/W-Subs) Long (WO/W-Subs)
-Proprietary Models
-GPT-4o (OpenAI, 2024) - $384^{2*}$ 71.2/77.8 80.1/82.2 71.3/76.8 62.2/74.3
-Gemini 1.5 Pro (Team et al., 2024) - 1/0.5 fps 75.0/81.3 81.7/84.5 74.3/81.0 67.4/77.4
-Open Source Model
-Video-LLaVA (Lin et al., 2023) 7B 8 39.9/41.6 45.3/46.1 38.0/40.7 36.2/38.1
-Long-LLaVA (Wang et al., 2024b) 7B 64 52.9/57.1 61.9/66.2 51.4/54.7 45.4/50.3
-mPLUG-Owl3 (Ye et al., 2024) 7B 128 59.3/68.1 70.0/72.8 57.7/66.9 50.1/64.5
-LongVA (Zhang et al., 2024b) 7B 128 52.6/54.3 61.1/61.1 50.4/53.6 46.2/47.6
-LongVILA (Xue et al., 2024) 7B 128 57.5/60.6 69.3/70.4 56.1/59.2 47.0/52.1
-256 57.4/61.8 69.0/70.8 56.7/60.6 46.4/54.0
-Train-Free Model or Key Frames selection
-Video-Tree* (Wang et al., 2024c) - - -/- -/- -/- 54.2/-
-LongVU* (Shen et al., 2024) 7B 1fps -/- -/- -/- -/60.6
-Chat-UniVi (Jin et al., 2024a) 7B 64 40.6/- 45.7/- 40.3/- 35.8/-
-MaxInfo Method
-Qwen2-VL-MaxInfo 7B Avg.180 62.1/70.0 72.5/76.4 62.0/69.8 51.8/63.9
-LLaVA-Video-MaxInfo 7B Avg.58 64.2/71.4 74.6/78.7 63.3/69.8 54.6/65.9
-A.3 Impact of MaxInfo Block’s Tol and Rank on Models.
-Overview of the details of how Tol and R affect the accuracy of the model, as Table 8, I.Frame is the number of frames input to the MaxInfo module, and F.Frame is the maximum number of frames after going through the MaxInfo Block.
-The effect of the initial number of samples on the model as shown in Table 9.
-Table 8: Comparison of experimental results on the effect of different tol and rank on the LLava-Video (Zhang et al., 2024c).
-Model Size R Tol (I.Frames, F.Frames) Acc.
-LLava-Video 7B 12 0.15 (96, 64) 60.35
-LLava-Video 7B 12 0.30 (96, 64) 61.25
-LLava-Video 7B 12 0.45 (96, 64) 61.10
-LLava-Video 7B 12 0.60 (96, 64) 61.14
-LLava-Video 7B 8 0.15 (96, 64) 60.06
-LLava-Video 7B 8 0.30 (96, 64) 61.48
-LLava-Video 7B 8 0.45 (96, 64) 59.68
-LLava-Video 7B 8 0.60 (96, 64) 57.80
-LLava-Video 7B 6 0.10 (96, 64) 61.10
-LLava-Video 7B 6 0.15 (96, 64) 60.96
-LLava-Video 7B 6 0.30 (96, 64) 60.58
-LLava-Video 7B 6 0.45 (96, 64) 59.01
-LLava-Video 7B 6 0.60 (96, 64) 57.89
-Table 9: Comparison of the effect of different initial sampling values on the model.
-Model Size (I.Frames, F.Frames) Acc.
-LLava-Video 7B (256, 64) 60.58
-LLava-Video 7B (192, 64) 60.73
-LLava-Video 7B (128, 64) 60.88
-LLava-Video 7B (96, 64) 60.06
+Such advancements have important implications for applications in education, accessibility, and public safety.
+However, improved video analysis capabilities may also raise ethical concerns, including potential misuse in surveillance, privacy violations, or biases affecting different communities.
+Ensuring responsible deployment with fairness and transparency is essential to mitigate these risks.
+In summary, while our approach provides significant benefits, its adoption should adhere to ethical principles to promote equitable and responsible use.